# Explore then Determine: A GNN-LLM Synergy Framework for Reasoning over Knowledge Graph

### v1

@@ 3.3 Knowledge-Enhanced Answer Determination
Although the exploration module can indicate the probabilities of candidates from the output scores, leveraging the powerful language understanding capabilities of the LLM for aligning the question linguistically with the explored information can provide greater benefits. Therefore, after exploring a set $\mathcal{C}_{q}$ of promising candidate answers with the relevant knowledge $\mathcal{K}_{q}$, we leverage LLM to determine the final answer. In this section, we design a knowledge-enhanced multiple-choice prompt to direct the LLM to generate final answers by leveraging both explicit knowledge from KG and implicit knowledge inside LLM.
Multiple-choice prompt. With the explored candidates $\mathcal{C}_{q}$, we design a multiple-choice format for LLM, which allows the LLM to consider and compare multiple given reference answers at once. The multiple-choice prompt is constructed using a task description, the provided question, and reference answers that contains the top-$N$ candidate entities in $\mathcal{C}_{q}$ identified with the highest probabilities during exploration. Meanwhile, we consider two kinds of knowledge aligned with each candidate to enhance the determination. First, we use the correct probability (as per Eq.(3)) of each reference answer to provide the confidence returned by the exploration module for LLM. Second, we extract the relevant path connecting from the topic entity $e_{q}$ to each candidate entity as an evidence chain.
These paths reveal the compositional associations between topic entities and candidate answers, providing fine-grained knowledge and faithful evidence for the final answer determination. Overall, the input prompt is formulated as on the right.
@@
Evidence extraction. To extract evidence chain as fine-grained compositional knowledge for LLM, we employ a greedy algorithm to trace back paths. Specifically, we backtrack from each reference candidate answer $e_{c}$, select the edge from $\widetilde{\mathcal{N}}_{e_{c}}^{L}$ with the highest attention weight (as per Eq.(1)), and set the head entity as a new starting point in the next step. By conducting $L$-steps backtracking, it will eventually trace back to the initial topic entity $e_{q}$ and obtain an evidence path between $e_{q}$ and $e_{c}$ (the detailed algorithm is shown in Appendix B.3). For example, in Figure 1, the evidence chain connecting the topic entity Birdy and the reference answer 1989 is: $\textit{Birdy}\xrightarrow{written\_by}\text{William Wharton}\xrightarrow{wrote}\text{Dad}\xrightarrow{release\_year}\textit{1989}$. Furthermore, LLMs can conduct semantic checks based on evidence chains to determine the final answer. For instance, it can recognize that the second reference answer (1976) in Figure 1 does not align with the semantic context of the question based on the paths between Birdy and 1976.
@@
Moreover, to avoid the LLM fully relying on the information given by the exploration module, and utilize the implicit knowledge within the LLM, we design the prompt of the task description in a way that encourages the LLM to generate answers based on its own knowledge. The complete input-output examples of the final answer determination are provided in the Appendix D, which demonstrates our method’s capability for accurate and interpretable reasoning. In addition, to avoid distractions from irrelevant information, we do not offer any few-shot examples, providing a clear structure for the LLM to follow, while maintaining the benefits of zero-shot learning.
@@ 3.4 Training Strategy
During training, we freeze the LLM in both text encoding and answer determination steps to avoid the expensive training cost. Instead, we only update the lightweight GNN and the scoring MLP in exploration module by minimizing the loss in Eq.(4). During reasoning, unlike traditional approaches [5, 7, 8], which first retrieve a subgraph and then perform reasoning on it, our exploration module can unify and simultaneously perform these two steps, adaptively filter out irrelevant information on the graph, and achieve more efficient reasoning. Additionally, it is noteworthy that the answer determination module can be adapted to any pre-trained LLM without the need for retraining or fine-tuning, thereby avoiding time and resource overhead.
Considering there are common compositional relationships between questions and the concepts in the graphs, the lightweight network in exploration module can be further benefited from pre-training strategies. Specifically, we pre-train the networks on two comprehensive KGQA datasets WebQSP [38] and CWQ [39], and then fine-tune them on target datasets. This approach enables the exploration module to better learn the compositional relationships in KGQA, enhancing question understanding and generalization abilities.
@@ 4.1 Experimental Setup
Experiment Details. In the pre-training stage, we set the dimension $d$ as 256 for the exploration module, learning rate as 1e-4, batch size as 20, number of layers $L$ as 3 and number of sampling $K$ as 200. As for the fine-tuning stage, we adjust the $L$ and $K$ based on the performance on validation set, and details are described in Appendix B.4. Considering the plug-and-play convenience of EtD, we use two LLMs for answer determination in experiments: Llama2-13B-chat [11], and ChatGPT [^1] https://openai.com/. We typically set number of reference answers $N$ as 3, and the influence of $N$ is shown in Appendix C.2. The pre-training and fine-tuning of exploration module are conducted on a RTX 3090-24GB GPU, and the inference of Llama2-13B-chat is running on two RTX 3090-24GB GPUs.
Baseline Methods. We consider the following baseline methods for performance comparison: (1) KG-based methods without using LLMs: KV-Mem [22], GraftNet [5], EmbedKGQA [23], NSM [7], SR+NSM [25], UniKGQA [8]; (2) LLM-based methods: KB-Binder [30] based on Codex [41], KAPING [15] based on GPT-3 [9], RoG [32] that can be plug-and-play with different LLMs, KD-CoT [35] based on ChatGPT, StructGPT [17] based on ChatGPT, ToG [18] that can be plug-and-play with different LLMs.
@@
GPT3-based KB-Binder 74.4 - 92.9 99.9 99.5
KAPING 73.9 - - - -
@@
EtD-ChatGPT 82.5 62.0 98.1 99.7 99.7
@@ 4.2 Performance Comparison
From the results in Table 2, it can be observed that our method EtD, whether combined with the Llama2-13B-chat or ChatGPT, outperforms traditional methods without LLMs, except for the MetaQA-2hop dataset, where many methods achieve over 99% accuracy. From the last two groups of the table, it can be seen that incorporating KG can effectively enhance the performance of LLMs, as they often lack the knowledge relevant to the questions. Within methods using Llama2, our approach combined with the 13B model outperforms ToG which is combined with the 70B model, and also outperforms some methods using ChatGPT. Similarly, within methods using ChatGPT, our approach demonstrates significant advantages. This shows the superiority of proposed explore-then-determine framework, effectively harnessing the precision of compositional learning and the prowess of language understanding. Additionally, it is evident that the performance of our method improves with the integration of more powerful LLM, e.g. ChatGPT. This demonstrates that our method effectively combines explicit structural knowledge from KGs with implicit knowledge from LLM, leading to more accurate reasoning and answering.
We also compare the inference time and number of interaction with LLM of several representative LLM-based methods, specifically executing Llama2-13B within our local environment. As can be seen in Table 3, RoG requires two interactions with LLM, while step-by-step reasoning methods StructGPT and ToG require even more frequent interactions, thus incurring a high cost. However, in our approach, using LLM for question encoding and graph exploration is highly efficient, and we ultimately requires only one-step inference of LLM, thereby achieving faster speeds.
4.3 Ablation Study
4.3.1 Effectiveness of Graph Exploration
In this section, we analyze the effectiveness of our exploration module, focusing on the Hits@1 of its independent output answer without the answer determination by LLM (i.e., EtD-w.o.-AD). Here, we choose UniKGQA, the state-of-the-art KG-based model without LLMs for comparison. As can be seen in Table 5, our exploration module outperforms UniKGQA on all datasets, demonstrating the effectiveness of the designed network for topology and semantic aware reasoning on the graph.
4.3.2 Effectiveness of LLM’s Answer Determination
Furthermore, Table 5 demonstrates the effectiveness of using LLM for answer determination. Although the exploration module itself (EtD-w.o.-AD) achieves decent performance, utilizing the powerful language processing capability of LLM for answer determination brings significant gains. With its powerful language processing capabilities, the LLM can analyze the semantic relationships between a given question, the candidate answers and the relevant knowledge. By leveraging the implicit knowledge embedded within itself, it achieves more accurate reasoning.
@@ 4.3.2 Effectiveness of LLM’s Answer Determination
EtD 1.29 1 1.99 1
Table 4: Comparison of different variants of EtD-Llama2-13B in Hits@1(%).
@@
EtD 77.4 57.7
w.o.-mcp 72.7 54.0
w.o.-cand 77.1 50.0
w.o.-prob 75.2 50.4
w.o.-path 76.8 57.0
Table 5: Comparison of Different Variants of EtD in Hits@1(%).
@@
EtD-w.o.-AD 76.4 55.6 97.6 99.1 99.5
EtD-Llama2-13B 77.4 57.7 97.9 99.1 99.6
EtD-ChatGPT 82.5 62.0 98.1 99.7 99.7
@@ 4.3.3 Influence of Prompt Form for Answer Determination
We design different forms of prompts to evaluate their influence in guiding the final answer determination of EtD-Llama2-13B. The variants include: (1) not using multiple-choice prompt (EtD-w.o.-mcp), (2) not using candidate entities (EtD-w.o.-cand), (3) not using correct probabilities (EtD-w.o.-prob), and (4) not using evidence paths (EtD-w.o.-path), which are shown in Appendix E.
As shown in Table 4, compared with EtD-w.o.-mcp, which inputs all candidate answers continuously without specific labels, our designed multiple-choice prompt can achieve better performance, since it offers a clearer structure and strength the connection of candidate answers with related knowledge. Moreover, all three factors in the prompt, i.e., candidate entity, correct probability and evidence path, have a positive impact on guiding LLM to make correct choices. Providing candidate answers can refine the target for LLM (compared with EtD-w.o.-cand), and correct probabilities prompts LLM to engage in reasoning with attention (compared with EtD-w.o.-prob). Evidence paths are crucial as they provide fine-grained knowledge for LLM reasoning (compared with EtD-w.o.-path), ensuring the transparency and reliability of the output results, as can be seen in Section 4.4.
4.3.4 Influence of Pruning
In this section, we discuss the influence of the pruning for candidate set. As can be seen in Figure 3, the size of the unfiltered candidate set and the number of edges involved in each question are typically immense due to the large-scale KG. In comparison, with pruning, both of them have significantly decreased. Therefore, we can filter out a large amount of irrelevant information, significantly reducing the computational costs, achieving more efficient reasoning.
(a) Number of candidates.
(b) Number of edges.
Figure 3: Influence of pruning on three datasets.
4.3.5 Effect of Pre-training Strategy
In Section 3.4, we have mentioned that the networks in exploration module can be benefited from pre-training strategies. We compare the performance of individual training strategy and pretrain-finetune strategy in Table 6. It is evident that the strategy of pre-training significantly outperforms training separately on distinct datasets. This underscores the effectiveness of pre-training in enabling our exploration module to better grasp the common compositional relationships between questions and the concepts in the graphs, thereby enhancing the question-comprehension abilities empowered by LLM. We also compare the impact of training strategy on the learning curves and conduct experiments on a new dataset in sports domain in the Appendix C.1.
Table 6: Influence of pretraining of exploration module.
Methods WebQSP CWQ MetaQA-1 MetaQA-2 MetaQA-3
w.o.-pretrain 72.6 53.4 97.8 99.0 99.4
pretrain-finetune 76.4 55.6 97.8 99.1 99.5
@@ 4.4 Case Study
We present two case studies in Figure 4, which display the question, reference answers of the input prompt, and the final result of the LLM’s answer determination. The shared task description can be found in Appendix D. It can be observed that the LLM makes correct determinations based on the candidate answers, correct probabilities, and relevant path facts, even when the right answer is not ranked at the top, indicating the importance of determination stage. Moreover, the LLM provides its reasoning process, enhancing the credibility of the output. These examples illustrate how our method combines explicit knowledge from the exploration stage with the implicit knowledge of the LLM, achieving accurate and faithful results.
@@
Figure 4: Examples of accurate and faithful output by EtD-Llama2-13B.
@@ 5 Conclusion
In this paper, we present a novel Explore-then-Determine (EtD) framework that synergizes LLMs with GNNs for reasoning over KG. During the Explore stage, we design an LLM-empowered GNN module to perform compositional learning on the KG, exploring the promising candidates and fine-grained compositional knowledge of the question. In the Determine stage, the information gathered from the exploration is organized to create a knowledge-enhanced multiple-choice prompt, which a frozen LLM then uses to determine the final answer. This two-step process leverages the precision of GNN-based compositional learning and the language understanding prowess of LLMs, effectively combining explicit knowledge from KGs with implicit knowledge in LLMs. Extensive experiments demonstrate the superiority of EtD, achieving accurate, efficient and faithful reasoning over KGs.

### v2

@@ 3.3 Knowledge-Enhanced Answer Determination
Although the GNN module can indicate the probabilities of candidates from the output scores, leveraging the powerful language understanding capabilities of the LLM for aligning the question linguistically with the explored information can provide greater benefits. Therefore, we extract high-quality reasoning chains to guide the LLM in reasoning for final answer determination. Specifically, to enhance the associative “System 1”-like reasoning of LLMs, we design a knowledge-enhanced multiple-choice prompt, guiding LLMs in effectively combining external explicit knowledge with its own internal implicit knowledge, and generating answers with a rapid one-step inference.
@@
Reasoning chain extraction. To extract valuable reasoning chains for LLM, we employ a greedy algorithm to trace back paths. Specifically, we first preserve top-$N$ candidate entities in $\mathcal{C}_{q}$ identified with the highest probabilities as reference answers. Then, we backtrack from each candidate answer $e_{c}$, select the edge from $\widetilde{\mathcal{N}}_{e_{c}}^{L}$ with the highest attention weight (as per (1)), and set the head entity as a new starting point in the next step. By conducting $L$-steps backtracking, it will eventually trace back to the initial topic entity $e_{q}$ and obtain a reasoning evidence chain between $e_{q}$ and $e_{c}$ (the detailed algorithm is shown in Appendix B.3). These chains reveal the compositional associations between topic entities and candidate answers, providing faithful evidence for the final answer determination. For example, in Figure 1, the evidence chain connecting the topic entity Birdy and the reference answer 1989 is: $\textit{Birdy}\xrightarrow{written\_by}\text{William Wharton}\xrightarrow{wrote}\text{Dad}\xrightarrow{release\_year}\textit{1989}$. In this way, LLMs can conduct semantic analysis and reasoning based on reasoning chains. For instance, it can recognize that the second reference answer (1976) in Figure 1 does not align with the semantic context of the question based on the chains between Birdy and 1976.
Knowledge-enhanced multiple-choice prompt. To effectively leverage the extracted reasoning chains, we further refine them into a knowledge-enhanced multiple-choice prompt. We incorporate two additional types of knowledge to enhance the LLM’s resoning for answer determination. First, we extract the list of candidate entity and assign them labels. This approach provides clear targets, effectively stimulating the LLM’s associative and intuitive reasoning capabilities. Second, we use the correct probability (as per (3)) of each candidate answer to provide the confidence returned by the GNN module for LLM, which aids the LLM in engaging in reasoning with heightened attention.
Overall, the knowledge-enhanced multiple-choice prompt is built using a task description, the provided question, and reference answers, which include candidate entity, correct probability and reasoning chain, as shown on the right.
@@
The complete input-output examples of the final answer determination are provided in the Appendix D, which demonstrates our method’s capability for accurate and interpretable reasoning. In addition, to avoid distractions from irrelevant information, we do not offer any few-shot examples, providing a clear structure for the LLM to follow, while maintaining the benefits of zero-shot learning.
@@ 3.4 Training Strategy
During training, we freeze the LLM in both text encoding and answer determination steps to avoid expensive costs. Instead, we only update the lightweight GNN module by minimizing the loss in (4). Moreover, unlike traditional GNN-based approaches [29, 30, 31], which first retrieve a subgraph and then perform reasoning on it, our GNN module can unify and simultaneously perform these two steps, adaptively filter out irrelevant information on the graph, and achieve more efficient reasoning. Additionally, it is noteworthy that the answer determination module can be adapted to any pre-trained LLM with just one-step inference, without the need for fine-tuning or frequent interaction, thereby avoiding time and resource overhead. In this way, the GNN module can served as “System 2”, as it performs explicit, step-by-step reasoning on graph structures, while the LLM aligns with “System 1”, leveraging implicit, associative patterns for rapid and intuitive decision-making.
Considering there are common compositional relationships between questions and the concepts in the graphs, the GNN module can be further benefited from pre-training strategies. Specifically, we pre-train the networks on two comprehensive KGQA datasets WebQSP [46] and CWQ [47], and then fine-tune them on target datasets. This approach enables the GNN module to better learn the compositional relationships, enhancing question understanding and generalization abilities.
@@ 4.1 Experimental Setup
GPT-3-based KB-Binder 74.4 - 92.9 99.9 99.5
KAPING 73.9 55.4 - - -
@@
DualR-ChatGPT 82.8 62.0 98.1 99.7 99.7
GPT-4-based GPT-4 67.3 46.0 65.7 34.6 48.9
ToG-GPT-4 82.6 69.5 - - -
DualR-GPT-4 87.6 73.6 98.3 99.9 99.9
Experiment Details. In the pre-training stage, we set the dimension $d$ as 256 for the GNN module, learning rate as 1e-4, batch size as 20, number of layers $L$ as 3 and number of sampling $K$ as 200. As for the fine-tuning stage, we adjust the $L$ and $K$ based on the performance on validation set, and details are described in Appendix B.4. Considering the plug-and-play convenience of DualR, we use three LLMs for answer determination in experiments: Llama2-13B-chat [3], ChatGPT and GPT-4111https://openai.com/. We typically set number of reference answers $N$ as 3, and the influence of $N$ is shown in Appendix C.4. The GNN module is trained on an RTX 3090-24GB GPU, while inference for Llama2-13B-chat runs on two RTX 3090-24GB GPUs.
Baseline Methods. We consider following baseline methods for performance comparison: (1) KG-based methods without using LLMs: KV-Mem [49], GraftNet [29], EmbedKGQA [28], NSM [30], SR+NSM [50], UniKGQA [31]; (2) LLM-based methods: KB-Binder [42] based on Codex [51], KAPING [15] based on GPT-3 [1], RoG [38] that can be plug-and-play with different LLMs, KD-CoT [37] and StructGPT [17] based on ChatGPT, ToG [18] that can be plug-and-play with different LLMs.
@@ 4.2 Performance Comparison
Main results. From the results in Table 2, it can be observed that our method DualR, whether combined with the Llama2 or GPT, outperforms traditional methods without LLMs. From the last three groups of the table, it can be seen that incorporating KG can effectively enhance the performance of LLMs, as they often lack the knowledge relevant to the questions. Within methods using Llama2, DualR combined with the 13B model outperforms ToG which is combined with the 70B model, and also outperforms some methods using ChatGPT. Similarly, within methods using ChatGPT and GPT-4, our approach demonstrates significant advantages. This shows the superiority of proposed dual-reasoning framework, effectively harnessing the GNN’s precise explicit learning and the LLM’s powerful language understanding, outperforming LLM’s single reasoning significantly. Additionally, the performance of DualR improves with the integration of more powerful LLM. This demonstrates that DualR effectively combines explicit knowledge from KGs with implicit knowledge from LLM, leading to more accurate reasoning.
Efficiency comparison. We also compare the inference time and number of interaction with LLM of several representative LLM-based methods, executing Llama2-13B within our local environment. As can be seen in Table 3, RoG requires two interactions with LLM, while step-by-step reasoning methods StructGPT and ToG require even more frequent interactions, thus incurring a high cost. However, in our approach, question encoding of LLM and graph reasoning of GNN are highly efficient, and we ultimately requires only one-step inference of LLM, thereby achieving high efficiency.
@@
DualR 1.29 1 1.99 1
Table 4: Comparison of different variants of DualR-Llama2-13B in Hits@1(%).
@@
DualR 78.3 58.0
w.o.-mcp 72.9 54.1
w.o.-cand 77.5 52.5
w.o.-prob 76.2 55.4
w.o.-chain 76.8 56.0
Fine-tuned setting results. While our method can seamlessly integrate with any off-the-shelf LLM, we expect that instruction fine-tuning the LLM for answer determination can lead to further performance improvements. We compare our approach with two recent state-of-the-art methods, RoG and GNN-RAG [39], both of which fine-tune Llama2-7B-chat on WebQSP and CWQ datasets, using Hit, Hits@1 and F1 as evaluation metrics. As shown in Table 5, DualR achieves state-of-the-art performance across multiple metrics, especially showing greater gains on more challenging dataset (i.e., CWQ), further validating the effectiveness of our method.
Table 5: Performance comparison of different methods with fine-tuned Llama2-7B-chat.
Methods WebQSP CWQ
Hit Hits@1 F1 Hit Hits@1 F1
RoG 85.7 80.0 70.8 62.6 57.8 56.2
GNN-RAG 85.7 80.6 71.3 66.8 61.7 59.4
DualR 84.9 81.5 71.6 68.9 65.3 62.1
4.3 Ablation Study
4.3.1 Effectiveness of Knowledge Exploration
In this section, we analyze the effectiveness of our GNN module, focusing on the Hits@1 of its independent output answer without the answer determination by LLM (i.e., DualR-w.o.-AD). Here, we choose UniKGQA, the state-of-the-art KG-based model without LLMs for comparison. As can be seen in Table 6, our GNN module outperforms UniKGQA on all datasets, demonstrating the effectiveness of the designed network for topology and semantic aware reasoning on the graph, capable of identifying promising candidates and extracting high-quality reasoning chains.
4.3.2 Effectiveness of LLM’s Answer Determination
Furthermore, Table 6 demonstrates the effectiveness of using LLM to reason for answer determination. Although the GNN module itself (DualR-w.o.-AD) achieves decent performance, leveraging extracted reasoning chains to guide LLM for answer determination brings significant gains. With its powerful language processing capabilities, the LLM can analyze the semantic relationships between a given question and relevant reasoning chains and make the thoughtful determination. By leveraging the implicit knowledge embedded within itself, it achieves more accurate reasoning.
Table 6: Comparison of Different Variants of DualR in Hits@1(%).
@@ 4.3.2 Effectiveness of LLM’s Answer Determination
DualR-w.o.-AD 76.8 55.6 97.6 99.1 99.5
DualR-Llama2-13B 78.3 58.0 97.9 99.1 99.6
DualR-ChatGPT 82.8 62.0 98.1 99.7 99.7
@@ 4.3.3 Influence of Prompt Form for Answer Determination
We design different forms of prompts to evaluate their influence in guiding answer determination of DualR-Llama2-13B. The variants include: (1) not using multiple-choice prompt (DualR-w.o.-mcp), (2) not using candidate entities (DualR-w.o.-cand), (3) not using correct probabilities (DualR-w.o.-prob), and (4) not using reasoning chains (DualR-w.o.-chain), which are shown in Appendix E. As shown in Table 4, compared with DualR-w.o.-mcp, which inputs all candidate answers continuously without specific labels, our designed multiple-choice prompt can achieve better performance, since it offers a clearer structure and strength the connection of candidate answers with reasoning chains. Moreover, all three factors in the prompt, i.e., candidate entity, correct probability and reasoning chain, have a positive impact on guiding LLM to make correct choices. Providing candidate answers can refine the target for LLM (compared with DualR-w.o.-cand), and correct probabilities prompts LLM to engage in reasoning with attention (compared with DualR-w.o.-prob). Reasoning paths are crucial as they provide detailed process for LLM reasoning (compared with DualR-w.o.-chain), enhancing the accuracy and reliability of the output, as can be seen in Section 4.4.
@@ 4.4 Case Study
We present two case studies in Figure 3, which display the question, reference answers of the input prompt, and the final result of the LLM’s answer determination. The task description can be found in Appendix D. It can be observed that the LLM without Dual-Reasoning (i.e., Llama2-13B-CoT) may suffer from hallucination and lack of knowledge, resulting in wrong answers. In contrast, DualR-Llama2-13B generates the correct answers, even when the right answer is not ranked at the top. Moreover, based on the given information, the LLM provides its reasoning process, which is consist with the input reasoning chains, enhancing the credibility of the output. These examples illustrate how DualR effectively implement LLM’s intuitive reasoning through externalized reasoning chains, achieving accurate and faithful results.
@@
Figure 3: Examples of accurate and faithful output by DualR-Llama2-13B.
@@ 5 Conclusion
In this paper, we introduce Dual-Reasoning (DualR), a novel framework following dual-process theory, which combines GNN-based structured, deliberate reasoning of “System 2” with LLM-based intuitive reasoning of “System 1” in KGQA task. DualR designs an LLM-empowered GNN module to explicitly reason on KGs, extracting high-quality reasoning chains, which are then refined into a knowledge-enhanced multiple-choice prompt, guiding the LLM to reason for final answer. With the collaboration of GNN and LLM, this two-tier reasoning process harnesses the precision graph learning of GNN and the language understanding prowess of LLM. Extensive experiments demonstrate the superiority of DualR, achieving accurate, efficient and faithful reasoning over KGs.
Acknowledgements
This work is supported by National Key Research and Development Program of China (under Grant No.2023YFB2903904), National Natural Science Foundation of China (under Grant No.92270106) and Beijing Natural Science Foundation (under Grant No.4242039).